OpenStudy (josee):

evaluate the integral I= integral between 0 and pi/2 of cosx{2f'(sinx)-3) dx when f(0)=3 and f(1)=5

OpenStudy (anonymous):

put sin(x)=t, cos(x)dx=dt I= \[\int\limits_{0}^{1} (2f'(t)-3) dt\] \[0[2f(t) - 3t]1\] =7-6 =1

OpenStudy (josee):

ohhhh. Thank you:)

OpenStudy (perl):

i dont understand the question here

OpenStudy (perl):